# references.bib

### v1

@@ Abstract
Recent robotic task planning frameworks have integrated large multimodal models (LMMs) such as GPT-4V. To address grounding issues of such models, it has been suggested to split the pipeline into perceptional state grounding and subsequent state-based planning. As we show in this work, the state grounding ability of LMM-based approaches is still limited by weaknesses in granular, structured, domain-specific scene understanding. To address this shortcoming, we develop a more structured state grounding framework that features a domain-conditioned scene graph as its scene representation. We show that such representation is actionable in nature as it is directly mappable to a symbolic state in classical planning languages such as PDDL. We provide an instantiation of our state grounding framework where the domain-conditioned scene graph generation is implemented with a lightweight vision-language approach that classifies domain-specific predicates on top of domain-relevant object detections. Evaluated across three domains, our approach achieves significantly higher state estimation accuracy and task planning success rates compared to the previous LMM-based approaches.
@@ I INTRODUCTION
Task planning in a real environment relies on two core capabilities: (a) reasoning to find an action plan that fulfills the goal, and (b) scene understanding to accurately recognize the state of the environment [sayplan]. Traditionally, these capabilities had to be learned through in-domain training, which resulted in models that could only perform well within specific tasks, objects, or environments. Large Language Models (LLMs) with their strong generalization offer the potential to overcome these limitations and have therefore gained considerable attention as general task planners [saycan]. Their extension to multimodality (LMMs) furthermore promised to enable joint reasoning over scene observation and instruction.
@@
The straightforward approach [replanvlm] would be to prompt the LMM with the observation and instruction inputs and let it directly output a task plan (Figure 1a). Studies [sayplan, groundeddecoding] pointed out that grounding them in reality is the main challenge. One point is they must adhere to robot- and domain-specific constraints such as the available skills, affordances and rules [sayplan]. The other side of grounding is they must identify the domain-relevant state of the environment, i.e. what are the relevant objects in the scene and what is their current configuration [neuroground]. Recent works [vilain, neuroground, vila, verigraph, grid] have therefore suggested to build on an intermediate state representation (Figure 1b.).
@@
Figure 1: a) Direct task planning from $I$mage, $D$omain specification and $G$oal instruction [replanvlm]. b) Enhancing grounding through an Intermediate Representation [vilain, neuroground, vila, verigraph]. c) We perform domain-conditioned scene graph generation (DC-SGG) to extract a representation in a structure that is directly mappable to plannable PDDL predicates.
@@
Among them, Shirai et al. [vilain] provide structure for constraints and state representation by utilizing Planning Domain Definition Language (PDDL)[pddl] which is used by classic symbolic planning systems. A PDDL domain specifies robot-specific skills and domain-specific lifted predicates, such as “carries” or “in”. The objective is then to transform input scene observation and task instruction into a PDDL problem, which grounds the predicates into initial state, e.g. “carries(robot,cucumber)”, and goal state, e.g. “in(cucumber,bowl)”. A symbolic planner can find an optimal plan that transforms initial state to goal state, even for large state spaces, a problem LLMs struggled at [llmp].
@@
Outsourcing (a) “finding an action plan” to a symbolic planner, shifts the problem focus to (b) “recognition of the environment state”. Previous works generate initial and goal state either by letting an LLM reason on top of detections of vision models [vilain], or, more recently, by directly processing scene observation, domain specification and task instruction in a single multimodal such as GPT-4V [vila, neuroground]. However, as we show in this paper, LMM-based parsing of the scene into the correct initial state is currently the major bottleneck for task planning success. We therefore aim for a more accurate scene parser that yields a plannable representation. We formulate the problem as domain-conditioned scene graph generation, where the task is to predict a set of task-relevant objects(nodes) and domain-relevant relationships(edges). This opens the door for leveraging ideas from scene graph research while maintaining a representation structure that complies with PDDL, and is hence, in contrast to general purpose scene graphs, directly actionable. We propose an LMM-free baseline that builds a simple relationship classifier on top of described [describedod] object detection. By comparing (Figure 1c.) our method with the different approaches for state grounding and state-based planning, we expose current limitations of LMM-based methods and the challenges to solve. Our contributions are therefore:
@@ II-A Task Planning from Observed State
Planning on top of a visually estimated state is a classic pipeline approach. Structuring the state representation in an object-centric and symbolic form [rpn, geometricsymbolic, sornet] further allows to find plans through logic search on the symbolic level. The symbolic state can be visually estimated through learning predicate classifiers [groundingpredicatesthroughactions, kase, sornet] grounded on object detections. However, these methods do no not generalize beyond a closed set of predicates and objects or environments seen during train-time. In contrast, due to their common world knowledge and reasoning abilities, large pre-trained models promise to overcome this limitation. Robotic task planning with large language models (LLMs) and vision language models (VLMs) has therefore been studied [sayplan, saycan, vilain], furthermore allowing the goal to be specified in natural language. Operating on text modality, visual grounded task planning with LLMs needs VLMs to transmit abstracted visual information to the LLM. This might lose task-relevant information and prohibits joint reasoning over vision and language modalities [vila]. Hence, multimodal LMMs [^1] We use the term LMM instead of VLM to distinguish approaches operating in open-ended text-generative LLM mode (LMM) from vision-centered models like open-vocabulary object detectors (VLM). came into the focus of recent research [replanvlm, vila, neuroground], of which [replanvlm] directly schedules actions, needing another LMM for error correction, [vila] builds a language intermediate scene representation through chain-of-thought prompting and [neuroground] explicitly prompts for initial state, goal state and action plan to be sequentially generated. In other words, these approaches use generated text as an intermediate representation which is not strictly bound to a structure.
@@ II-B PDDL Problem as State Representation
Tasks that should follow more structured, well-defined and human-interpretable [rpn] constraints can be specified in a formal language such as Planning Domain Definition Language (PDDL) [pddl]. A PDDL domain file sets the constraints of robot and domain, whereas a PDDL problem file consists of the set of present objects, initial state and goal state. Kase et al. [kase] learn to predict the PDDL initial state from observation, but requires in-domain training, hence limited to seen objects, predicates and actions. Research on LLM-planning provided the state representation in PDDL [generalizedplanning], or equipped the LLM with an algorithmic PDDL solver to achieve correct optimal plans [llmp]. The latter is similar to our motivation, we use a PDDL solver and follow the problem setting of [vilain] that the domain file specifying the constraints is given and the problem file specifying the scene’s initial and goal state is to be generated.
@@ II-C Scene Graph as a Plannable Scene Representation
Ray et al. [tampin3dsg] perform Task and Motion Planning (TAMP) in given hierarchical 3D scene graphs designed for large scale environments. While they also use PDDL predicates, they need to infer them on top of the predefined graph structure, which restricts the set of inferrable predicates and thus also the range of solvable tasks. Also within a TAMP framework, Zhu et al. [geometricsymbolic] construct scene graphs at the geometric and symbolic level, allowing motion planning on the geometric and task planning on the symbolic scene graph. Their basic idea to map a geometric scene graph to a symbolic scene graph is similar to ours, but the mapping is manually defined and the set of predicates is fixed. ConceptGraphs [conceptgraphs] incrementally build a 3D scene graph through fusing the predictions of vision models into a set of entities and annotating their properties and relations with the help of vision-language models. The shortcoming of this approach is that the decomposition of the scene and generation of annotations is generic and not task- and domain-conditioned, hence structure and information may not align with the target task. Among the LLM-based works [sayplan, llmstate, grid, verigraph], SayPlan [sayplan] lets an LLM search and plan in a 3D scene graph, solving scalability issues. Ni et al. [grid] transform scene graphs to tokens to train an LLM-Encoder Action-Decoder network, showing superior performance over LLM-as-Planner approaches, but requiring training on scene- and robot-specific data. VeriGraph [verigraph] and [sequentialmanipulationonsg] model actions as graph edit operations, whereof [verigraph] uses LMMs for both scene graph generation and planning on the scene graph. Similar to other LMM-based methods [sayplan, replanvlm, vilain], the open-ended generative nature of LLMs leads to incorrect propositions, requiring multiple iterations of error-corrective reprompting.
@@ IV METHOD
$f_{state}$ from the problem definition needs to generate initial and goal state. We identified that generating the correct initial state is the main weakness in previous frameworks. Therefore, in Section IV-A we dedicate our attention to the design of a scene parser. The goal parser is then covered by Section IV-B. Unlike previous frameworks, we generate the goal state first, independent from the harder to estimate initial state, avoiding error propagation. Given initial and goal states and benefiting from the structured plannable representation, we can follow [vilain] to use an off-the-shelf solver [fastdownward] for $f_{plan}$.
@@ IV-A1 Object Detection
We use a described object detector to find the set of domain- and task-relevant objects $\mathcal{O}$. Described object detection [describedod] unifies open-vocabulary object detection and phrase grounding, where the former detects all instances that match a text query class (e.g. cucumber), and the latter detects the single instance which matches the referring text phrase (e.g. the cucumber to the right of the bowl). We need the former for detecting all objects that are domain-relevant, i.e. included in types $D_{t}$, and the latter for uniquely grounding the instances that are mentioned in the goal specification/instruction $G$. Hence, the object detector is queried on image $I$ with types $D_{t}$ as class queries and phrases extracted from $G$ as phrase grounding queries. The output consists of two sets of bounding boxes, one based on $D_{t}$, the other based on $G$. We merge the detections by matching each box based on $G$ with one box based on $D_{t}$ by maximum Intersection over Union (IoU) to obtain the set of $N$ vertices $V$ for the spatial scene graph:
@@
where each vertex has a type, name and box. PDDL Objects can be written from this: $\{(n_{i},t_{i})\}_{i=1}^{N},$ e.g. cucumber - vegetable or white_bowl - container.
@@ IV-A2 Spatial to Semantic Scene Graph
The goal is now to map this spatial scene graph to semantic scene graph $(V,E)$, with edges $E$ built from domain-relevant predicates $D_{p}$. We achieve this through conditioning on the domain’s set of types $D_{t}$ and predicate functions $D_{p}$, using $D_{p}$’s type constraint on its object arguments. For example, for a binary predicate like “in”, the first object argument can be of general “object” type while the second must be of “container” type. We leverage this constraint for the scene graph and consider only triplets that may form a valid predicate. For each predicate $p\in D_{p}$, the set of valid edges(triplets) is built for both an example (one-shot) and the test image, resulting in $E_{p}^{test},E_{p}^{ex}$. Their corresponding spatial relations $R_{p}^{s,test},R_{p}^{s,ex}$ are considered features for classifying semantic relationships $D_{p}$. Say a triplet $e=(v_{i},p,v_{j})$ evaluates $True$ iff. $p(v_{i},v_{j})$ holds. Then each test triplet $e\in E_{p}^{test}$ gets classified by evaluating the predicate of its nearest neighbor (NN) triplet in $E_{p}^{ex}$. NN is calculated from Euclidean distance in the feature space shared by $R_{p}^{s,test},R_{p}^{s,ex}$. Finally, all edges that are classified as $True$ are included in scene graph $\mathcal{G}=(V,E)$:
@@ V-A Metrics
Additionally, for checking the generated PDDL, we adopt the measure of problem file validity and plan validity222determining validity with VAL (https://github.com/KCL-Planning/VAL) from [vilain] and present results in Table III.
@@ V-B Dataset
Our evaluation is based on the Problem Description Generation dataset ProDG [vilain]. It provides three domains ($D$), each defined by a PDDL domain file specifying object types, lifted predicates and lifted actions. In each domain there is a set of problems to solve. Each problem comes with an image ($I$) as observation and a natural language instruction ($G$).
TABLE I: ProDG-v characteristics. Observed Predicates, number of derived predicates, #A number of skills in skillset, mmA median minimum actions (that are required to complete the task), m#O median number of objects.
@@
TABLE II: State grounding results (middle) in Precision|Recall of scene graph triplets. Task planning results (right) in success rate. OV denotes open-vocabulary, S2S denotes our spatial to semantic classifier. IR lists type of intermediate representation, T text, $\mathcal{G}$ scene graph, (:p PDDL predicates.
@@
VILA[vila] 0 LMM GPT-4o T - - - GPT-4o 0.00 0.10 0.20
NeuroGround[neuroground] 1 LMM GPT-4o T 1.00|0.92 0.74|0.86 0.41|0.50 GPT-4o 1.00 0.37 0.08
@@
VeriGraph[verigraph] 1 LMM GPT-4o $\mathcal{G}$ 0.92|0.92 0.60|0.85 0.39|0.32 GPT-4 - - -
ViLaIn[vilain] 1 VLM →LLM GDINO +BLIP2 →GPT-4 (:p 0.79|0.72 0.49|0.42 0.85|0.72 Solver 0.00 0.03 0.10
[…]
@@
[…]
To better suit the state grounding problem setting, we modify the original ProDG to create ProDG-v(isual). This is necessary because the original ProDG dataset included (i) predicates that cannot be determined from observation, their truth value would need to be guessed based on few-shot examples, (ii) predicates that are always true across the dataset, diluting metrics. Hence, in ProDG-v we distinguish between observed and derived predicates, where observed predicates can be visually perceived whereas derived predicates are merely a consequence of observed predicates or the consequence(effect) of an action. Observed predicates can then be predicted based on the image and derived predicates logically inferred based on PDDL definitions within the domain file.

### v2

# Domain-Conditioned Scene Graphs for State-Grounded Task Planning

@@ Abstract
Recent robotic task planning frameworks have integrated large multimodal models (LMMs) such as GPT-4o. To address grounding issues of such models, it has been suggested to split the pipeline into perceptional state grounding and subsequent state-based planning. As we show in this work, the state grounding ability of LMM-based approaches is still limited by weaknesses in granular, structured, domain-specific scene understanding. To address this shortcoming, we develop a more structured state grounding framework that features a domain-conditioned scene graph as its scene representation. We show that such representation is actionable in nature as it is directly mappable to a symbolic state in planning languages such as the Planning Domain Definition Language (PDDL). We provide an instantiation of our state grounding framework where the domain-conditioned scene graph generation is implemented with a lightweight vision-language approach that classifies domain-specific predicates on top of domain-relevant object detections. Evaluated across three domains, our approach achieves significantly higher state grounding accuracy and task planning success rates compared to LMM-based approaches. https://github.com/Vision-Kek/DC-SGG
@@ I INTRODUCTION
Task planning in a real environment relies on two core capabilities: (a) reasoning to find an action plan that fulfills the goal, and (b) scene understanding to accurately recognize the state of the environment [1]. Traditionally, these capabilities had to be learned through in-domain training, which resulted in models that could only perform well within specific tasks, objects, or environments. Large Language Models (LLMs) with their strong generalization offer the potential to overcome these limitations and have therefore gained considerable attention as general task planners [2]. Their extension to multimodality (LMMs) furthermore promised to enable joint reasoning over scene observation and instruction.
@@
The straightforward approach [3] would be to prompt the LMM with the observation and instruction inputs and let it directly output a task plan (Figure 1a). Studies [1, 4] pointed out that grounding them in reality is the main challenge. One point of grounding is to enforce adherence to robot- and domain-specific constraints such as the available skills, affordances and rules [1]. The other side of grounding is they must identify the domain-relevant state of the environment, i.e. what are the relevant objects in the scene and what is their current configuration [5]. Recent works [6, 5, 7, 8, 9] have therefore suggested to build on an intermediate state representation (Figure 1b).
@@
Figure 1: a) Direct task planning from IImage, DDomain specification and GGoal instruction [3]. b) Enhancing grounding through an Intermediate Representation [6, 5, 7, 8]. c) We perform domain-conditioned scene graph generation (DC-SGG) to extract a representation in a structure that is directly mappable to plannable PDDL predicates.
@@
Among them, Shirai et al. [6] provide structure for constraints and state representation by utilizing Planning Domain Definition Language (PDDL) [10] which is used by classic symbolic planning systems. A PDDL domain specifies robot-specific skills and domain-specific lifted predicates, such as “carries” or “in”. The objective is then to transform input scene observation and task instruction into a PDDL problem, which grounds the predicates into initial state, e.g. “carries(robot,cucumber)”, and goal state, e.g. “in(cucumber,bowl)”. A symbolic planner can find an optimal plan that transforms initial state to goal state, even for large state spaces, a problem LLMs struggled at [11].
@@
Outsourcing (a) “finding an action plan” to a symbolic planner, shifts the problem focus to (b) “recognition of the environment state”. Previous works generate initial and goal state either by letting an LLM reason on top of detections of vision models [6], or, more recently, by directly processing scene observation, domain specification and task instruction in a single multimodal such as GPT-4o [7, 5]. However, as we show in this paper, LMM-based parsing of the scene into the correct initial state is currently the major bottleneck for task planning success. We therefore aim for a more accurate scene parser that yields a plannable representation. We formulate the problem as domain-conditioned scene graph generation, where the task is to predict a set of task-relevant objects (nodes) and domain-relevant relationships (edges). This opens the door for leveraging ideas from scene graph research while maintaining a representation structure that complies with PDDL, and is hence, in contrast to general purpose scene graphs, directly actionable. We propose an LMM-free baseline that builds a simple relationship classifier on top of described [12] object detection. By comparing (Figure 1c) our method with the different approaches for state grounding and state-based planning, we expose current limitations of LMM-based methods and the challenges to solve. Our contributions are therefore:
@@ II-A Task Planning from Observed State
Planning on top of a visually estimated state is a classic pipeline approach. Structuring the state representation in an object-centric and symbolic form [13, 14, 15] further allows to find plans through logic search on the symbolic level. The symbolic state can be visually estimated through learning predicate classifiers [16, 17, 15] grounded on object detections. However, these methods do no not generalize beyond a closed set of predicates and objects or environments seen during train-time. In contrast, due to their common world knowledge and reasoning abilities, large pre-trained models promise to overcome this limitation. Robotic task planning with large language models (LLMs) and vision language models (VLMs) has therefore been studied [1, 2, 6], furthermore allowing the goal to be specified in natural language. Operating on text modality, visual grounded task planning with LLMs needs VLMs to transmit abstracted visual information to the LLM. This might lose task-relevant information and prohibits joint reasoning over vision and language modalities [7]. Hence, multimodal LMMs [^1] We use the term LMM instead of VLM to distinguish approaches operating in open-ended text-generative LLM mode (LMM) from vision-centered models like open-vocabulary object detectors (VLM). came into the focus of recent research [3, 7, 5], of which [3] directly schedules actions, needing another LMM for error correction, [7] builds a language intermediate scene representation through chain-of-thought prompting and [5] explicitly prompts for initial state, goal state and action plan to be sequentially generated. In other words, these approaches use generated text as an intermediate representation which is not strictly bound to a structure.
@@ II-B PDDL Problem as State Representation
Tasks that should follow more structured, well-defined and human-interpretable [13] constraints can be specified in a formal language such as Planning Domain Definition Language (PDDL) [10]. Kase et al. [17] learn to predict the PDDL initial state from observation, but require in-domain training, hence are limited to seen objects, predicates and actions. Research on LLM-planning provided the state representation in PDDL [18], or equipped the LLM with an algorithmic PDDL solver to achieve correct optimal plans [11]. The latter is similar to our motivation, we use a PDDL solver and follow the problem setting of [6] that the domain file specifying the constraints is given and the problem file specifying the scene’s initial and goal state is to be generated.
@@ II-C Scene Graph as a Plannable Scene Representation
Ray et al. [19] perform Task and Motion Planning (TAMP) in given hierarchical 3D scene graphs designed for large scale environments. While they also use PDDL predicates, they need to infer them on top of the predefined graph structure, which restricts the set of inferrable predicates and thus also the range of solvable tasks. Also within a TAMP framework, Zhu et al. [14] construct scene graphs at the geometric and symbolic level, allowing motion planning on the geometric and task planning on the symbolic scene graph. Their basic idea to map a geometric scene graph to a symbolic scene graph is similar to ours, but the mapping is manually defined and the set of predicates is fixed. ConceptGraphs [20] incrementally build a 3D scene graph through fusing the predictions of vision models into a set of entities and annotating their properties and relations with the help of vision-language models. The shortcoming of this approach is that the decomposition of the scene and generation of annotations is generic and not task- and domain-conditioned, hence structure and information may not align with the target task. Among the LLM-based works [1, 21, 9, 8], SayPlan [1] lets an LLM search and plan in a 3D scene graph, solving scalability issues. Ni et al. [9] transform scene graphs to tokens to train an LLM-Encoder Action-Decoder network, showing superior performance over LLM-as-Planner approaches, but requiring training on scene- and robot-specific data. VeriGraph [8] and [22] model actions as graph edit operations, whereof [8] uses LMMs for both scene graph generation and planning on the scene graph. Similar to other LMM-based methods [1, 3, 6], the open-ended generative nature of LLMs leads to incorrect propositions, requiring multiple iterations of error-corrective reprompting.
@@ IV METHOD
$f_{state}$ from the problem definition needs to generate initial and goal state. We identified that generating the correct initial state is the main weakness in previous frameworks. Therefore, in Section IV-A we dedicate our attention to the design of a scene parser. The goal parser is then covered by Section IV-B. Unlike previous frameworks, we generate the goal state first, independent from the harder to estimate initial state, avoiding error propagation. Given initial and goal states and benefiting from the structured plannable representation, we can follow [6] to use an off-the-shelf solver [23] for $f_{plan}$.
@@ IV-A1 Object Detection
We use a described object detector to find the set of domain- and task-relevant objects $\mathcal{O}$. Described object detection [12] unifies open-vocabulary object detection and phrase grounding, where the former detects all instances that match a text query class (e.g. cucumber), and the latter detects the single instance which matches the referring text phrase (e.g. the cucumber to the right of the bowl). We need the former for detecting all objects that are domain-relevant, i.e. included in types $D_{t}$, and the latter for uniquely grounding the instances that are mentioned in the goal specification/instruction $G$. Hence, the object detector is queried on image $I$ with types $D_{t}$ as class queries and phrases extracted from $G$ as phrase grounding queries. The output consists of two sets of bounding boxes, one based on $D_{t}$, the other based on $G$. We merge the detections by matching each box based on $G$ with one box based on $D_{t}$ by maximum Intersection over Union (IoU) to obtain the set of $N$ vertices $V$ for the spatial scene graph:
@@
where each vertex has a type $t_{i}$, name $n_{i}$ and box $b_{i}$. PDDL Objects can be written from this: $\{(n_{i},t_{i})\}_{i=1}^{N},$ e.g. cucumber-vegetable or white_bowl-container.
@@ IV-A2 Spatial to Semantic Scene Graph
The goal is now to map this spatial scene graph to semantic scene graph $(V,E)$, with edges $E$ built from domain-relevant predicates $D_{p}$. We achieve this through conditioning on the domain’s set of types $D_{t}$ and predicate functions $D_{p}$, using $D_{p}$’s type constraint on its object arguments. For example, for a binary predicate like “in”, the first object argument can be of general “object” type while the second must be of “container” type. We leverage this constraint for the scene graph and consider only triplets that may form a valid predicate. For each predicate $p\in D_{p}$, the set of valid edges (triplets) is built for both an example (one-shot) and the test image, resulting in $E_{p}^{test},E_{p}^{ex}$. Their corresponding spatial relations $R_{p}^{s,test},R_{p}^{s,ex}$ are considered features for classifying semantic relationships $D_{p}$. Say a triplet $e=(v_{i},p,v_{j})$ evaluates $True$ iff. $p(v_{i},v_{j})$ holds. Then each test triplet $e\in E_{p}^{test}$ gets classified by evaluating the predicate of its nearest neighbor (NN) triplet in $E_{p}^{ex}$. NN is calculated from Euclidean distance in the spatial feature space shared by $R_{p}^{s,test},R_{p}^{s,ex}$. Finally, all edges that are classified as $True$ are included in scene graph $\mathcal{G}=(V,E)$:
@@ V-A Metrics
Additionally, for checking the generated PDDL, we adopt the measure of problem file validity and plan validity from [6] and present results in Table III.
@@ V-B Dataset
Our evaluation is based on the Problem Description Generation dataset ProDG [6]. It provides three domains ($D$), each defined by a PDDL domain file specifying object types, lifted predicates and lifted actions. In each domain there is a set of problems to solve. Each problem comes with an image ($I$) as observation and a natural language instruction ($G$).
TABLE I: ProDG-v characteristics: Observed Predicates, number of derived predicates, #A: number of skills in skillset, mmA: median minimum actions (that are required to complete the task), m#O: median number of objects.
@@
TABLE II: State grounding results in Precision|Recall of scene graph triplets. Task planning results in success rate. OV denotes open-vocabulary, S2S denotes our spatial to semantic classifier. IR lists type of intermediate representation, specifically T text, $\mathcal{G}$ scene graph, and (:p PDDL predicates.
@@
VILA[7] 0 LMM GPT-4o T - - - GPT-4o 0.00 0.10 0.20
NeuroGround[5] 1 LMM GPT-4o T 1.00|0.92 0.74|0.86 0.41|0.50 GPT-4o 1.00 0.37 0.08
@@
VeriGraph[8] 1 LMM GPT-4o $\mathcal{G}$ 0.92|0.92 0.60|0.85 0.39|0.32 GPT-4 - - -
ViLaIn[6] 1 VLM →LLM GDINO +BLIP2 →GPT-4 (:p 0.79|0.72 0.49|0.42 0.85|0.72 Solver 0.00 0.03 0.10
Ours 1 $\text{DC-SGG}_{VQA}$ GDINO →GPT-4o $\mathcal{G}$(:p 1.00|1.00 0.71|0.79 0.41|0.36 Solver 1.00 0.32 0.24
@@
Ours 1 $\text{DC-SGG}_{S2S}$ GDINO→S2S $\mathcal{G}$(:p 1.00|1.00 0.98|1.00 0.94|0.91 Solver 0.94 0.86 0.65
To better suit the state grounding problem setting, we modify the original ProDG to create ProDG-v(isual). This is necessary because the original ProDG dataset included (i) predicates that cannot be determined from observation, their truth value would need to be guessed based on few-shot examples, (ii) predicates that are always true across the dataset, diluting metrics. Hence, in ProDG-v we distinguish between observed and derived predicates, where observed predicates can be visually perceived whereas derived predicates are merely a consequence of observed predicates or the consequence(effect) of an action. Observed predicates can then be predicted from the image, while derived predicates are logically inferred based on domain definitions.